Find the area of the paraboloid x2 + y2 = z inside the cylinder x2 + y2 = 9.


1
Expert's answer
2021-12-07T05:00:27-0500

"z_x = 2x"

"z_y=2y"

The surface area over the region R defined by "x^2+y^2 = 9=3^2" is

"S = \int \int_R \sqrt{(z_x)^2+(z_y)^2+1}dxdy"

= "\int \int_R \sqrt{4x^2+4y^2+1}dxdy"

Then to polar coordinates

"S = \int_{0}^{2\pi} \int_0^3(4r^2+1)^{1/2}rdrd\theta"

"= \dfrac{1}{12} \int_{0}^{2\pi} (4r^2+1)^{3/2}|_0^3d\theta"

"= \dfrac{37\sqrt{37} -1}{12} \int_0^{2\pi} d\theta"

"=\dfrac{37\sqrt{37} -1}{12} 2\pi"

"= \pi \dfrac{37\sqrt{37}-1}{6}"